The sum of two numbers is -52. The first number minus the second is -30 . find the numbers
Answer by mananth(16949)

let x, y be the numbers
x+y =-52
x-y=-30
..
add the two equations
2x=-82
/2
x= -41
..
x+y = -52
plug x
-41+y=-52
y=-52+41
y=-11
...
m.ananth@hotmail.ca